OpenStudy (anonymous):

state the slope and y intercept for x+y=5

OpenStudy (acxbox22):

change the equation into a slope intercept form by subtracting x from both sides y=-x+5 we now have it in y=mx+b m is the slope and b is the y intercept

OpenStudy (anonymous):

But I don't understand how to get the answers...?

OpenStudy (acxbox22):

slope is the number in front of the x and y intercept is the number added in this case the slope is -1 and the y intercept is (0,5)

OpenStudy (anonymous):

Oh, okay. Thanks!